OpenStudy (anonymous):

The asymptotes of q(x) = 1/2 csc 2x are

OpenStudy (anonymous):

\[q(x)=1/2\sin2x\] asymptote denominator is zero \[2\sin 2x=0\] \[\sin 2x=0\] \[2x=2.\pi.k,k \epsilon \mathbb{R}\] \[x=2.\pi.k\]

OpenStudy (anonymous):

\[\csc x=1/\sin x\]

OpenStudy (anonymous):

so x is \[(\pm 2\pi,\pm4\pi...)\]